OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x + 2y = -6 3x + y = 8 A. -2y + 6 B. 2y - 6 C. 2y + 6 D. -2y - 6

OpenStudy (rogue):

Well, isolating x in the first equation means that we have to get it alone on one side of the equation. -x + 2y = -6 We can first subtract 2y from both sides. -x + 2y - 2y = -6 - 2y -x = -6 - 2y Next, we just multiply both sides by negative 1 to make the x positive. -x*-1 = (-6 - 2y)*-1 x = 6 + 2y = 2y + 6 Choice C.